Consider a population consisting the scores of 6 students in a Statistic Test.




18, 22, 25, 28, 32, 36




Suppose samples of size 3 are drawn from this population, describe the sampling




distribution of the sample means following the steps below.





Step 4: List all possible




samples and their




corresponding means.

1
Expert's answer
2022-04-01T14:02:44-0400

"\textnormal{Sample\qquad\qquad Mean}\\ 
\ \\
18,22,25\qquad \cfrac{18+22+25}{3}=21\frac{2}{3}\\
18,22,28\qquad \cfrac{18+22+28}{3}=22\frac{2}{3}\\
18,22,32\qquad \cfrac{18+22+32}{3}=24\\
18,22,36\qquad \qquad 25\frac{1}{3}\\
18,25,28\qquad \qquad 23\frac{2}{3}\\
18,25,32\qquad \qquad 25\\
18,25,36\qquad \qquad 26\frac{1}{3}\\
18,28,32\qquad \qquad 26\\
18,28,36\qquad \qquad 27\frac{1}{3}\\
18,32,36\qquad \qquad 28\frac{2}{3}\\
22,25,28\qquad \qquad 25\\
22,25,32\qquad \qquad 26\frac{1}{3}\\
22,25,36\qquad \qquad 27\frac{2}{3}\\
22,28,32\qquad \qquad 27\frac{1}{3}\\
22,28,36\qquad \qquad 28\frac{2}{3}\\
22,32,36\qquad \qquad 30\\
25,28,32\qquad \qquad 28\frac{1}{3}\\
25,28,36\qquad \qquad 29\frac{2}{3}\\
25,32,36\qquad \qquad 31\\
28,32,36\qquad \qquad 32."